### v2

# Sa2VA: Marrying SAM2 with LLaVA for Dense Grounded Understanding of Images and Videos

Abstract
This work presents Sa2VA, the first unified model for dense grounded understanding of both images and videos. Unlike existing multi-modal large language models, which are often limited to specific modalities and tasks, Sa2VA supports a wide range of image and video tasks, including referring segmentation and conversation, with minimal one-shot instruction tuning. Sa2VA combines SAM-2, a foundation video segmentation model, with LLaVA, an advanced vision-language model, and unifies text, image, and video into a shared LLM token space. Using the LLM, Sa2VA generates instruction tokens that guide SAM-2 in producing precise masks, enabling a grounded, multi-modal understanding of both static and dynamic visual content. Additionally, we introduce Ref-SAV, an auto-labeled dataset containing over 72k object expressions in complex video scenes, designed to boost model performance. We also manually validate 2k video objects in the Ref-SAV datasets to benchmark referring video object segmentation in complex environments. Experiments show that Sa2VA achieves state-of-the-art across multiple tasks, particularly in referring video object segmentation, highlighting its potential for complex real-world applications. Code, dataset, and models are released for further research.
Figure 1: Illustration of capabilities of our proposed Sa2VA. (a). Given a video, Sa2VA is able to segment the referred object and understand the whole scene. (b).Sa2VA supports image conversation, video conversation, image referring segmentation, video referring segmentation, and grounded caption generation with single-shot instruction-tuning. (c).Sa2VA achieves strong results on multiple images, video referring segmentation, and chat benchmarks compared with existing MLLMs, such as GLaMM [66] and OMG-LLaVA [99].
1 Introduction
Multi-modal Large Language Models (MLLMs) have made significant progress with the rapid development of Large Language Models (LLMs) [73, 33, 86]. They have benefited from various image and video-level tasks such as visual question answering (VQA) [1, 70], narrative story generation [87, 26, 105], and interactive editing [76, 39, 30].
One important direction is to understand video content in a fine-grained manner, including segmenting and tracking pixels with language descriptions, and performing VQA on visual prompts in the video. In particular, we want to achieve promptable fine-gained analysis of video since the user can be in the loop when playing the video in an interactive mode, as shown in Fig. 1 (a). This leads to various applications, such as short video editing [6, 75, 58], robot navigation [23, 24, 21], and surveillance analysis [63].
However, either state-of-the-art video perception models [11, 48, 98, 67] or video MLLMs [95, 50, 55, 43] cannot achieve this. The formers are limited by the constrained semantic concepts, lacking the open-ended ability (Video VQA, or named video chat tasks). For example, the latest foundation model, SAM-2 [67], can achieve promptable segmentation and tracking. However, it cannot perform text-aware tasks, such as understanding language expressions or video conversation. On the other hand, video MLLMs [51, 55, 42, 50] can understand long videos and perform VQA. For example, the latest LLaVA [42] can achieve strong results on video VQA. However, it cannot carry out perception tasks, and also it cannot understand visual prompts. Several works [99, 18, 85] explore the combination of perception models and MMLMs. However, most works explore image datasets or try to solve one specific video understanding task. To our knowledge, no work can leverage the benefits from both sides.
Through the above analysis, in this work, we make the first step to unify two visual foundation models, SAM-2 [67] and LLaVA-like MLLMs [53], into one framework. In this way, our model not only inherits the spatial-temporal perception capabilities of SAM-2 and the open-ended abilities of MLLMs, but also benefits from learning additional knowledge derived from their fine-grained training data. However, to achieve this goal, three key challenges must be addressed: (1) Task formulation: How to effectively formulate a range of tasks in a one-shot training setting, particularly for multimodal inputs. In particular for video perception tasks and video VQA tasks. Previous works [85, 18] only solve partial of them. (2) Performance balance: How to resolve conflicts between tasks, such as ensuring strong referring visual understanding abilities without sacrificing the language proficiency of MLLMs. Note that previous works [99, 82] find conversation tasks degraded when performing grounding tasks. (3) Knowledge sharing: How to leverage the pre-trained knowledge from SAM-2 and MLLMs and build a robust, unified model. SAM-2 is trained with more than 1B masks, and most MLLMs are trained with massive instructing pairs.
We present Sa2VA, the first model that integrates SAM-2 with LLaVA-like MLLMs, creating a unified, grounded understanding of both images and videos. We formulate various tasks into a one-shot visual instruction tuning format, including referring segmentation, visual question answering (VQA), and grounded conversation generation (GCG) for both image and video. Our key insight is to leverage the flexible token-length handling of LLMs, treating all input images, videos, and visual prompts as visual tokens without requiring additional specific design. Through joint co-training, we demonstrate that conflicts between grounding and conversation tasks can be effectively resolved. Unlike existing approaches that use MLLMs as agents or connectors to call on visual experts, our model is trained end-to-end, showcasing scalability in both the model and the datasets. Moreover, we adopt a decoupled design in which SAM-2’s decoder and memory module are frozen, allowing us to retain the perception and tracking capabilities of SAM-2. This design also makes our method a plug-and-play module, enabling our model to be updated with the latest MLLMs.
In addition, we introduce a challenging referring video segmentation dataset, Ref-SAV, curated through an automatic pipeline based on the source SA-V datasets [67]. This is from our empirical observation that existing referring video segmentation datasets are limited to small-scale and short clips with fewer occlusions. We benchmark several existing models on this dataset and find ample room for improvement and further exploration.
Table 1: Comparison of capabilities of different representative models. Our method supports various tasks and modalities. Benefiting from these interactive features on video, Sa2VA can perform multiple promptable tasks in the video, as shown in Fig. 1 (a) and (b).
Method Support Inputs Dense Grounding Conversation End to End
Image Video Visual Prompts RES Ref-VOS Inter-VOS GCG Image-Chat Video-Chat Video Caption Interactive Caption Training
[…]
Sa2VA is co-trained with multi-dataset co-training, including image and video datasets. Without any bells and whistles, Sa2VA achieves state-of-the-art results on six referring image and video segmentation datasets while keeping strong image and video chat ability compared with previous MLLMs, as shown in Fig. 1 (c). On Ref-SAV, our method achieves over 15% previous methods under zero-shot test setting and further stronger results using our proposed Ref-SAV training set. As shown in Fig. 1 (c), our work builds a new strong baseline for a unified, dense, grounded understanding of images and video. Overall, our contributions are as follows:
  • From the visual token perspective, we re-think several tasks, including image chat, image referring segmentation, video chat, referring video object segmentation, and grounded caption generation. We formulate all these tasks as a single instruction-tuning process.
  • We develop Sa2VA, the first simple framework that combines SAM-2 and LLaVA-like models into one model, where we achieve spatial-temporal grounded understanding. We introduce a simple decouple framework with a shared SEG token. We also present a simple mask tracking method by re-utilizing SAM-2’s knowledge.
  • We introduce a challenging referring video object segmentation benchmark, Ref-SAV, with a manual check for each example. The benchmark introduces heavier occlusions, long text inputs, motion blurs. We also develop a simple data annotation pipeline to build the Ref-SAV training dataset, where we find the training dataset improve model performance on Ref-SAV.
2 Related Work
Multi-modal Large Language Models. Earlier works [44, 45, 31] explore better multi-modal fusion methods and feature extractors and design various fusion architectures, particularly for vision language tasks. With the advanced LLMs frameworks [4, 73, 33], multi-modal instruction tuning on LLMs [2, 71, 53, 8] becomes the main pipeline. Then, various benchmarks [32, 59, 49, 19, 72] come up and serve as better sources to achieve stronger performance, meaning that data plays a central role in current MLLMs. One representative work, LLaVA [53], treats the visual features as visual tokens. The authors also provide a unified data format to unify extensive VQA tasks. After that, several works [92] explore stronger visual cues to enhance the visual inputs of LLaVA. Meanwhile, several works [96, 94, 68, 97, 14, 15, 51, 95, 65, 29, 40] add extra components to adapt LLaVA for visual grounding, detection, segmentation, and video VQA. Recently, there has also been a new trend [42, 55] to unify image, video, and multi-image analysis in one framework. LLaVA-OneVision [42] design a single model to handle four different input sources. In visual perception, several works [84, 48, 67] also explore multi-dataset and multi-task co-training. SAM-2 [67] designs a unified system for joint image and video interactive segmentation. Our model, Sa2VA, integrates SAM-2 into current VLM models into one end-to-end model and aims to unify image and video for dense, grounded understanding, including segmentation, chat, and caption.
Referring Segmentation. This task aims to output specific masks (for image inputs) or tracked masks (for video input) driven by language description. Earlier works [91, 52, 38, 13, 80, 102] explore various fusion modules to achieve better performance. Then, several stronger models [84] adopt transformer-based methods to achieve unified segmentation and tracking in video. Equipped with LLMs, several recent works [40, 99, 83, 65] propose more complex referring tasks, including reasoning referring or joint mask and caption generation. For example, LISA [40] involves reasoning-based segmentation. Then, GLaMM [66] annotates a new dataset and proposes region-level caption and segmentation tasks. Meanwhile, several recent works explore joint instruction tuning on referring segmentation and conversation [85, 101, 8]. Our method extends these studies into the video domain with the help of SAM-2 [67], while keeping stronger performance on both image/video referring tasks and VQA tasks.
Video Segmentation and Grounding. Current video segmentation methods [47, 34, 106] focus on segmenting and tracking pixels in closed sets. Several works [25, 104] explore open-vocabulary settings. However, the concepts are still limited compared with the knowledge space of LLMs. For video grounding, recent works [28] explore LLM into the joint understanding of video and audio. Recent work VISA [85] also explores reasoning video object segmentation. However, it is trained solely on video segmentation datasets and lacks end-to-end training, which limits its capability in other tasks and the potential of scaling up. Our model, Sa2VA, supports fine-grained spatial-temporal modeling of both static (image) and dynamic (video) visual contents and achieves strong performance on various tasks.
3 Method
Figure 2: Our proposed Sa2VA model. The model first encodes the input texts, visual prompts, images, and videos into token embeddings. These tokens are then processed through a large language model (LLM). The output text tokens are used to generate the “[SEG]” token and associated language outputs. The SAM-2 decoder receives the image and video features from the SAM-2 encoder, along with the “[SEG]” token, to generate corresponding image and video masks.
3.1 Unifying Multi-task Representations
Developing a unified model for addressing diverse image and video understanding tasks is challenging due to the distinct formats of these tasks. To overcome this, we begin by reexamining the task formulations and propose a unified representation approach, laying the groundwork for the development of Sa2VA.
Referring Image/Video Object Segmentation. For image referring segmentation, given input text tokens $T_{i}\in\mathbb{R}^{N\times D}$, the model takes input images $I_{i}\in\mathbb{R}^{H\times{W}\times 3}$ and outputs corresponding binary masks $M_{o}\in\mathbb{R}^{H\times{W}}$ that align with text description. $N$ and $D$ are the number and dimension of text tokens. For video object referring segmentation, the model takes input videos $V_{i}\in\mathbb{R}^{T\times H\times{W}\times 3}$ and outputs binary spatio-temporal masks (masklets) $M_{o}\in\mathbb{R}^{T\times H\times{W}}$. $T$, $H$, and $W$ are video frame number, height, and width.
Image/Video Chat and Grounded Caption Generation. For image and video chat tasks, given input text tokens $T_{i}$ and corresponding images $I_{i}$ or video $V_{i}$, the model outputs answer text $T_{o}$. For grounded caption generation tasks, the model jointly outputs the masks $M_{o}$ and aligned text $T_{o}$.
Visual Prompt Understanding Tasks. For visual prompt understanding tasks, in addition to text tokens $T_{i}$ and image $I_{i}$, the model further takes extra visual prompt tokens $VP_{i}$ (boxes or points on the image) as inputs and output corresponding masks $M_{o}$ and aligned text answers $T_{o}$.
Unified Task Representation. Existing works solve the above tasks via specific design models or partial unified models (different tasks have different model weights). In this work, we point out that all the above tasks can be unified as a one-shot instruction-tuning process, since we can leverage the LLMs to handle various visual tokens. The overall process can be formulated as the following:
$T_{o},M_{o}=LLM(\{I_{i},V_{i},VP_{i}\},T_{i}).$ (1)
For chat-only tasks, the model only outputs text tokens $T_{o}$. For referring segmentation tasks, the model outputs masks (image) or masklets (video) $M_{o}$. For grounded caption generation tasks, the model outputs both texts and masks. For different tasks, the input visual tokens are changing according to specific tasks. Since $M_{o}$ (whether image masks or masklets) can be controlled with SEG tokens, one SEG token represents one prompt input of SAM-2’s decoder. Thus, it is natural to put all these tasks into one shared encoder-and-decoder framework and perform end-to-end training with one LLM and the SAM decoder.
Figure 3: Data annotation pipeline. Our proposed automatic data annotation pipeline consists of three stages: object-level, scene-level, and video-level text expression annotation. We use different colors in the final expression to highlight the information derived from each stage. Best view on screen and zoom in.
3.2 Sa2VA Framework
The overall architecture of Sa2VA is shown in Fig. 2. It contains two parts: the LLaVA-like model and SAM-2.
Pre-trained MLLMs. We adopt pre-trained LLaVA-like models as the MLLMs. It contains one visual encoder, one visual projection layer, and one LLM. The visual encoder takes input images, video, and sub-images as inputs. The visual projection layer maps inputs into visual tokens. These tokens, combined with the input text tokens, are the input of LLMs and the LLMs generate the text token prediction based on them. Note that we adopt pre-trained MLLMs following previous works [99, 66, 40, 85] to leverage their strong capability. For both image and video chat datasets, we follow the same pipeline [9, 2] without modification.
Decoupled Design. We append SAM-2 alongside the pre-trained LLaVA model. We do not take the SAM-2’s output tokens (visual features or decoder outputs) into LLM. There are three reasons. First, we want to make the combination as simple as possible without increasing extra computation costs. Secondly, adding extra tokens needs an extra alignment process. Thirdly, via this design, we can fully make our work as a plug-in-play framework to utilize pre-trained MLLMs since the MLLM community goes fast. Thus, we adopt a decoupled design without introducing further communication between LLaVA and SAM-2.
Tuning SAM-2 Decoder via SEG Tokens. We connect SAM-2 and the MLLM via the special token “[SEG]”. The hidden states of the “[SEG]” token are used as a new type of prompt and fed into SAM-2’s Decoder, where they are decoded into segmentation masks. The hidden states of “[SEG]” can be seen as a novel spatial-temporal prompt for SAM-2. SAM-2 segments the corresponding object mask in image and video based on the spatial-temporal prompt. During training, the SAM-2 decoder can be tuned to understand the spatial-temporal prompt, and gradients can be backpropagated through the “[SEG]” token to the MLLM, allowing it to output the spatial-temporal prompt better.
Utilizing SAM-2 knowledge for mask tracking. For Ref-VOS tasks, we design a simple framework to achieve strong results on public benchmarks. In particular, for giving input video, we adopt a “[SEG]” token to generate the masks of the key frames. Then, we use the memory encoded by the key frame features to generate the mask for the remaining frames. We detailed this process in the appendix.
3.3 Ref-SAV Dataset and Benchmark
Data Annotation Pipeline. We have carefully designed an automatic annotation pipeline to generate referring object text expressions for the SA-V dataset [67]. As shown in Fig. 3, the pipeline consists of 3 stages:
  1. 1.
    Object-level annotation. We first select the frame with the largest object area from the video and mask out the non-object pixels. The cropped and full images are then fed separately into InternVL2-76B [9] to generate detailed descriptions. The descriptions are processed by Qwen2-72B [86] for consistency checking, with conflicting descriptions discarded.
  2. 2.
    Scene-level annotation. Yellow contours are used to highlight the object in the image. Both the image and the object-level description from the previous stage are sent into InternVL2-76B [9] to generate a detailed object description that includes relationships to the scene and surrounding objects.
  3. 3.
    Video-level annotation. We uniformly sample 8 frames from the video, applying yellow borders to highlight the object in each frame. These frames, along with the scene-level description, are processed by InternVL2-76B to generate a video-level description, capturing the object’s movement and action.
Table 2: Experiment results on image/video referring segmentation benchmarks and image/video chat benchmarks. For MME dataset, A / B means the perception (A) and cognition (B) scores, while scores without “/” mean the total score (A + B)
Method Image Segmentation Video Segmentation Image Chat Video Chat GCG
RefCOCO [36] RefCOCO+ [36] RefCOCOg [90] MeViS [13] Ref-DAVIS17 [38] Ref-YTVOS [69] ReVOS [85] MME [19] MMBench [59] SEED-Bench [41] Video-MME [20] MMBench-Video [17] GCG [66]
LLAVA-1.5-13B [54] - - - - - - - 1531 68.8 70.1 - - -
Video-LLaVA-7B [51] - - - - - - - - 60.9 - 39.9 1.03 -
LLaMA-VID-7B [50] - - - - - - - 1521 65.1 59.9 - 1.08 -
mPLUG-Owl3-8B [89] - - - - - - - - 77.6 - 53.5 1.35 -
InternVL2-8B [9] - - - - - - - - 81.7 76.2 54.0 1.28 -
PixelLM-7B [68] 73.0 66.3 69.3 - - - - 309/135 17.4 - - - -
LaSagnA [77] 76.8 66.4 70.6 - - - - 0/0 0.0 - - - -
LISA-7B [40] 74.1 62.4 66.4 - - - - 1/1 0.4 - - - -
GLaMM-7B [66] 79.5 72.6 74.2 - - - - 14/9 36.8 - - - 28.9
LLaVA-G-7B [96] 77.1 68.8 71.5 - - - - - - - - - -
GSVA-13B [83] 79.2 70.3 75.7 - - - - - - - - - -
OMG-LLaVA-7B [99] 78.0 69.1 72.9 - - - - 1177/235 47.9 56.5 - - 29.9
VISA-13B [85] 72.4 59.8 65.5 44.5 70.4 63.0 50.9 - - - - - -
Sa2VA-1B (Ours) 77.4 69.9 72.3 41.7 72.3 65.3 47.6 1381/405 68.3 64.8 39.9 1.07 23.8
Sa2VA-4B (Ours) 78.9 71.7 74.1 46.2 73.8 70.0 53.2 1536/530 77.3 73.3 50.4 1.23 28.2
Sa2VA-8B (Ours) 81.6 76.2 78.7 46.9 75.2 70.7 57.6 1617/511 81.6 75.1 52.1 1.34 31.0
Sa2VA-26B (Ours) 82.5 78.8 79.7 46.2 77.0 70.1 58.4 1691/538 83.7 76.8 52.6 1.45 33.5
Ref-SAV Training Dataset. Using the above pipeline, we have automatically annotated detailed object expressions for the SA-V dataset, resulting in the creation of Ref-SAV. This dataset contains 37,311 videos and 72,509 object expressions. Notably, we refrain from augmenting the object expressions, ensuring that each object has only one corresponding expression.
Ref-SAV Evaluation Benchmark. We select a subset of videos from the SA-V dataset’s training set to construct the Ref-SAV evaluation benchmark, as the validation and test sets contain only a limited number of videos. These videos are strictly orthogonal to Ref-SAV’s training dataset. The evaluation benchmark consists of two parts: 1) The long-expression set, generated using the above automatic annotation pipeline and carefully filtered by human annotators. 2) The short expression set, which is entirely manually annotated. The evaluation benchmark includes 1,147 videos and 1,945 object expressions, of which 1,694 are long expressions and 251 are short expressions. For training and testing referring examples, please refer to the appendix.
3.4 Sa2VA Training and Testing.
One For All Co-Training. Sa2VA is co-trained multi-datasets. For VQA tasks, we adopt text regression loss $L_{text}$ as common MLLMs. For segmentation tasks $L_{mask}$, we apply pixel-wised cross-entropy loss $L_{CE}$ and dice loss $L_{DICE}$. Not that there is no pre-trained stag as previous works [53], we perform supervised fine-tuning in one-shot training.
$\mathcal{L}_{intruction}=\mathcal{L}_{text}+\mathcal{L}_{mask},\quad\mathcal{L}_{mask}=\mathcal{L}_{CE}+\mathcal{L}_{DICE}.$ (2)
In Sec. 4, in addition to co-training for one model, we further report specific tuned models on single tasks following  [66, 99] to enable fair comparison.
One For All Testing. All tasks can be encompassed within the Eq. 1 paradigm. During the inference stage, we encode the necessary task requirements, such as text prompts, visual prompts, image, and video features, into tokens to input into the LLM. The output tokens of LLM are then decoded into text responses (LLM prediction head), segmentation mask (SAM-2 decoder), and SAM-2 mask tracking module responses according to the task definition. We strongly recommend that the readers check the detailed model inference table in the appendix.
4 Experiments
Baseline. We construct the baseline by combining the SOTA MLLM models like InternVL2 [9] and SAM2 [67]. Similar to previous works  [40, 66, 99], the segmentation mask is obtained by decoding the hidden state of the “[SEG]” token through SAM2’s decoder. Inspired by Mask2Former-VIS [10], an object shares the same “[SEG]” token throughout the video frames, allowing our model to handle both image and video referring segmentation tasks in a unified way. In addition, we further use more advanced MLLM models, InternVL2-5 to show the model scaling effect of Sa2VA.
Table 3: Datasets used for experiments.
Type Datasets
Image QA LLaVA 1.5 (665K)
Image Segmentation RefCOCO (17K), RefCOCO+ (17K), RefCOCOg (22K), Grand-f (214K)
Video QA ChatUniVi (100K)
Video Segmentation Ref-YTVOS (3.5K), MeVIS (0.6K), ReVOS (1.7K), Ref-SAV (37K)
Table 4: Performance on image-level benchmarks with MLLMs with segmentation capability. Our method achieves the best accuracy trade-off between image chat and referring segmentation datasets.
Method MME [19] MMBench [59] SEED-Bench [41] AI2D [37] MMStar [7] MMMU [93] SQAtest [60] RefCOCO RefCOCO+ RefCOCOg
LISA-7B [40] 1/1 0.4 - 0.0 - - - 74.1 62.4 66.4
PixelLM-7B [68] 309/135 17.4 - 0.0 - - - 73.0 66.3 69.3
LaSagnA-7B [77] 0/0 0.0 - 0.0 - - - 76.8 66.4 70.6
GLaMM-7B [66] 14/9 36.8 - 28.2 - - - 79.5 72.6 74.2
OMG-LLaVA-7B [99] 1177/235 47.9 56.5 42.9 - - - 78.0 69.1 72.9
Sa2VA-4B (ours) 1553/540 76.8 72.6 79.9 53.7 46.2 95.8 80.4 74.3 76.7
Sa2VA-8B (ours) 1651/578 82.4 75.5 82.1 60.3 44.7 96.8 81.9 76.5 78.9
Table 5: Ref SAV validation sets. zs: zero-shot testing. ft: trained with our proposed Ref-SAV training dataset.
Method Long Short Overall
J F J&F J F J&F J F J&F
UniRef++ [79] (zs) 14.1 10.8 12.5 9.0 8.2 8.6 11.6 9.5 10.5
UNINEXT [84] (zs) 11.7 8.3 10.0 5.8 4.4 5.1 8.8 6.4 7.6
MeVIS [13] (zs) 12.1 7.1 11.3 6.2 5.3 5.5 12.2 9.8 10.3
VISA [85] (zs) 16.1 12.2 14.1 12.3 9.6 9.2 13.2 11.3 11.8
Sa2VA-8b (zs) 47.7 50.9 49.3 31.5 35.0 33.3 39.6 43.0 41.3
Sa2VA-8b (ft ) 57.0 60.4 58.7 39.5 42.9 41.2 48.3 51.7 50.0
Datasets and Metrics. We use four types of datasets to train Sa2VA, including image QA, video QA, image segmentation, and video segmentation datasets. As shown in Tab. 3, Sa2VA’s training data contains approximately 1.1M image-text or video-text pairs. Since InternVL2 has been trained with a large amount of image QA and video QA data, we only used 665K LLaVA 1.5 [57] and 100K ChatUniVi [35] data to prevent the MLLM from forgetting its image and video QA capabilities. We used 56K referring expression data [36, 90] and 214K grounding conversation generation data [66] for image-level text-driven segmentation. For video-level referring expression segmentation, we used 5.8K existing referring VOS data from Ref-YouTubeVOS [69], MeVIS [13], and ReVOS [85]. Additionally, we used 37K long text referring VOS data generated by our proposed automatic annotation pipeline to enhance Sa2VA’s understanding of long referring text and its object grounding capabilities for complex videos. For image referring segmentation, we adopt cIoU. For referring video object segmentation, we adopt J&F. For image and video chat tasks, we follow previous works [53, 50] and report performance accordingly.
Implementation Details. We adopt XTuner [12] codebase for training and testing. During the instruction tuning stage, the initial learning rate is set to 4e-5, with only the perception model kept frozen, and the LLM is fine-tuned with LoRA [27]. The maximum sequence length in the LLM is set to 8,192. All training is conducted on eight NVIDIA A800 GPUs with 80GB of memory. The instruction tuning stage took 48 hours. We adopt VLMEvalKit [16] for chat evaluation. For our benchmark testing, we adopt original open-sourced codebase [84, 79, 85] and model weights to inference the video results. A more detailed setup for each task can be found in the appendix.
4.1 Main Results
Comparison With SOTA MLLMs. As shown in Tab. 2, Sa2V-8B achieves 81.6, 76.2, and 78.9 cIoU on RefCOCO, RefCOCO+, and RefCOCOg, respectively, surpassing GLaMM-7B by 2.1, 3.6, and 4.5 cIoU. Undoubtedly, Sa2VA sets new SOTA results on RefCOCO+ and RefCOCOg, significantly outperforming previous grounding MLLMs, including LISA, GLaMM, PixelLLM, PSALM, and OMG-LLaVA. Meanwhile, Sa2VA also demonstrates strong conversational capabilities, achieving 2128, 81.6, and 75.1 scores on MME, MMbench, and SEED-Bench, respectively, while existing grounding MLLMs perform poorly in conversation. Sa2VA achieves comparable performance to InternVL2 on image QA benchmarks, demonstrating that Sa2VA largely preserves the base MLLM InternVL2’s image chat performance.
Sa2VA also performs excellently on video benchmarks. Sa2VA achieves 57.0, 75.2, and 57.6 J&F on MeVIS, Ref-DAVIS17, and ReVOS, respectively, surpassing the previous SOTA VISA-13B by 12.5, 4.8, and 6.7 J&F. Furthermore, Sa2VA-8B achieves a score of 1.34 on the video QA benchmark MMBench-Video, even surpassing InternVL2-8B’s 1.28. The main experimental results demonstrate that our Sa2VA is a versatile and powerful MLLM. Due to page limitations, more results can be found in the appendix.
Table 6: Comparison with Fine-tuned Models.
Model Type RefCOCO RefCOCO+ RefCOCOg
LAVT [88] 72.7 62.1 61.2
GlaMM-7B [66] 79.5 72.6 74.2
OMG-LLaVA-7B [99] 78.0 69.1 72.9
F-LLM-7B [82] 76.1 65.2 68.5
Sa2VA-4B (ours) 80.4 74.3 75.7
Sa2VA-8B (ours) 82.3 77.3 79.3
Detailed Results on Image-Level Benchmarks. As indicated by  [99, 40, 82], performing instruction tuning hurts the performance on chat tasks. We take co-training to handle this problem. In Tab. 4, we further report more image chat datasets. Our model can still achieve strong results on multiple image chat benchmarks while keeping SOTA referring segmentation results.
Results on Ref-SAV validation set. In Tab. 5, we benchmark several state-of-the-art Ref-VOS models using our proposed Ref-SAV benchmark. We find that even the foundation model [84] and recent video MLLM model [85] cannot achieve strong results on our benchmarks. This is because our benchmark has more occlusion, longer text description, and part-whole annotations from SAM-2 [67]. Conversely, our method, Sa2VA, even with open-sourced datasets (Tab. 3), can achieve strong results. Our method can be boosted further with our training set, which means that our Ref-SAV training set is a good supplementary to the video understanding community.
Comparison with specific fine-tuned methods. Following previous works [66, 99], we also compare con-current MLLMs in specific fine-tuned datasets. As shown in Tab. 6, we find 1.5% improvements on three datasets over the co-training model, and Sa2VA works best among recent works. It should be noted that our model with a smaller size can still achieve stronger results than previous MLLMs.
4.2 Ablation Studies
Table 7: Ablation study on co-training effect on multiple datasets.
Data Image Segmentation Video Segmentation Image Chat Video Chat
RefCOCO RefCOCO+ RefCOCOg MeViS Ref-DAVIS17 MME MMBench SEED-Bench Video-MME MMBench-Video
All Data 77.4 69.9 72.3 50.8 72.3 1381/405 68.3 64.8 39.9 1.07
w/o Image QA 78.0 70.1 72.2 48.3 73.0 1298/359 63.4 63.8 39.7 0.39
w/o Image Segmentation 20.2 20.6 23.2 38.0 48.8 1393/408 70.1 65.7 41.2 1.08
w/o Video QA 78.0 70.4 72.6 50.7 74.3 1370/402 69.1 65.0 41.3 0.71
w/o Video Segmentation 77.4 69.1 72.4 44.4 69.0 1403/398 67.8 64.9 40.4 1.04
Table 8: Experiment results using stronger InternVL2.5 in our Sa2VA.
Base Image Segmentation Video Segmentation Image Chat
MLLM RefCOCO [36] RefCOCO+ [36] RefCOCOg [90] MeViS (val_u) [13] MeViS (val) [13] Ref-YTVOS [69] Ref-DAVIS17 [38] MME [19] MMBench [59] SEED-Bench [41] AI2D [37] MMStar [7] SQAtest [60]
InternVL2.0-4B 80.4 74.3 76.7 52.1 46.2 70.0 73.8 1553/540 76.8 72.6 79.9 53.7 95.8
InternVL2.0-8B 81.9 76.5 78.9 57.0 46.9 70.7 75.2 1651/578 82.4 75.5 82.1 60.3 96.8
InternVL2.5-1B 79.6 73.6 77.7 53.4 47.0 68.0 69.5 1504/434 71.9 71.0 69.2 48.6 89.9
InternVL2.5-4B 82.4 77.6 79.7 55.9 46.4 71.3 73.7 1691/610 81.8 74.9 81.4 57.9 96.8
InternVL2.5-8B 82.6 78.0 80.3 58.9 51.5 72.3 75.9 1690/610 84.4 76.5 82.7 62.4 97.4
InternVL2.5-26B 82.9 79.3 81.2 61.8 52.1 75.1 78.6 1698/653 85.8 78.3 85.7 67.0 98.4
Effectiveness of Joint Co-training. Sa2VA’s superior performance on image and video QA and segmentation benchmarks is attributed to unified instruction tuning and joint co-training. We conducted ablation studies on the training datasets, and the results are shown in Tab. 7. When training without image QA datasets, Sa2VA’s performance on MME and MMBench drops by 129 and 4.9 scores, respectively. When training without image segmentation datasets, Sa2VA only achieves 20.2, 20.6, and 23.2 cIoU on RefCOCO, RefCOCO+, and RefCOCOg, while video segmentation benchmark performance declined significantly. Sa2VA’s performance on MMBench-Video decreases by 34% when training without video QA datasets. When training without video segmentation datasets, Sa2VA’s performance on MeVIS and Ref-DAVIS17 drops by 4.4 and 3.3 J&F, respectively. The ablation results demonstrate that joint co-training with these four types of datasets is key to Sa2VA’s exceptional performance across the various tasks.
Ablation study on the Segmentation Token Design. We have explored several segmentation token designs, with the results shown in Tab. 9. When using the repeat strategy, which makes the LLM output N “[SEG]” tokens for N frames, this approach slightly decreases video segmentation performance due to being more prone to missing or repeating “[SEG]” tokens. We also explored using different segmentation tokens for different frames, such as setting “[SEG 1]” to “[SEG n]” to segment the object across frames 1 to n. However, this prevented knowledge sharing with image segmentation tasks, resulting in a larger performance drop.
Scale ability of Sa2VA’s Data. In the main experiments, we have demonstrated that our Sa2VA can scale with model size. We conduct ablation experiments on additional data to verify that it can scale with data, as shown in Tab. 10. When we add 3M image QA data from Infinity-MM [22] (only stage-4), Sa2VA-1B shows a 2.1 score improvement on MMBench, with minimal impact on the image and video segmentation benchmark performance. Using our annotated Ref-SAV data, Sa2VA-1B exhibits a 1.7 J&F performance boost on the MeVIS benchmark. These results indicate that Sa2VA’s performance can continue to improve as the scale of SFT data increases.
Figure 4: The samples of our Ref-SAV benchmark. Our proposed benchmark features multi-granularity, complex occlusion and reappearing, and both short and long-format text expressions.
Visualization examples. We provide some demos of Sa2VA, shown in Fig. 1. Additionally, in Fig. 4, we showcase the features of our Ref-SAV benchmark through samples that demonstrate multi-granularity, occlusion reappearance, and both short and long text expressions. These features make Ref-SAV a challenging and comprehensive referring VOS benchmark. We provide more visualization results, ablation studies, and analysis in the appendix. We strongly recommend readers to check that.
Table 9: Ablation study on “[SEG]” token design.
Type RefCOCO RefCOCO+ RefCOCOg DAVIS MeVIS
Single 77.4 69.9 72.3 72.3 50.8
Repeat 77.3 70.2 72.5 71.1 49.6
Multiple 77.6 70.3 72.4 68.6 46.3
Table 10: Ablation study on using more datasets.
Dataset Size RefCOCO RefCOCOg MMBench MME MeVIS
baseline 1.2M 77.4 72.3 68.3 1381/405 50.8
Inifinity-MM [22] 1.2M+3M 77.1(-0.3) 72.6(+0.3) 70.4(+2.1) 1396/346(-44) 51.2(+0.4)
Ref-SAV 1.2M+37K 77.2(-0.2) 72.6(+0.3) 68.2(-0.1) 1384/418(+16) 52.5(+1.7)
5 Conclusion
In this work, we present Sa2VA, a versatile framework that integrates SAM-2, a foundation video segmentation model, with LLaVA-like MLLMs, to achieve a dense, grounded understanding of both images and video. Our method can handle various image and video understanding tasks, including referring image/video segmentation and image/video conversation, with just one-shot instruction tuning. By leveraging the knowledge from both LLaVA and SAM-2, our model has strong capabilities in both mask and language generation. To verify the effectiveness of our proposed method, we propose a challenging referring video object segmentation benchmark, Ref-SAV. Based on our experimental results, Sa2VA achieves strong performance on various benchmarks from different tasks.
6 Appendix
Overview. Our supplementary includes the following aspects:
  • Appendix A gives more experiment results on our Sa2VA models, datasets effectiveness and more tasks.
  • Appendix B presents more detailed descriptions of the model, a comparison with existing MLLMs, and inference pipelines.
  • Appendix C shows more detailed process of training dataset generation process.
  • Appendix D gives visual examples of Sa2VA on various dense grounded tasks.
  • Appendix E discuss the shortcomings of Sa2VA, future directions of our Ref-SAV benchmarks, and board impacts.
Appendix A More Experiment Results
Table 11: Comparison with Recent Video MLLMs.
Model Type MeViS ReVOS Ref-DAVIS17
PG-Video-LLaVA [61] 18.87 - -
GLaMM + SAM2 [62] 38.7 - -
VideoGLaMM [62] 45.2 - -
ViLLa [103] - - 64.4
VISA-13B [85] 44.5 50.9 70.4
VideoLISA-3.8B [3] 44.4 - 68.8
Sa2VA-4B (ours) 52.1 53.2 73.8
Table 12: Comparison with Vision Expert Models.
Model Type RefCOCO RefCOCO+ RefCOCOg MeVIS Ref-DAVIS17
LAVT [88] 72.7 62.1 61.2 - -
ReferFormer [78] - - - 31.0 61.1
UniRef++-L [79] 81.4 74.0 76.0 - 67.2
EVF-SAM [100] 82.4 76.5 78.2 - -
LMPM [13] - - - 37.2 -
UniVS [46] - - - - 59.4
Sa2VA-26B (ours) 82.5 78.8 79.7 57.3 77.0
Table 13: Sa2VA with Qwen2-VL model.
MLLM Type RefCOCO RefCOCO+ RefCOCOg Ref-DAVIS17 MeVIS ReVOS
Sa2VA: Qwen2-VL-2B [2] 76.9 68.7 72.9 72.0 49.4 40.0
Sa2VA: Intern-VL2-1B 77.4 69.9 72.3 72.3 50.8 47.6
Sa2VA: Intern-VL2-4B 78.9 71.7 74.1 73.8 52.1 53.2
Table 14: Ref SAV validation sets. ft: trained with our proposed Ref-SAV training dataset. zs: zero-shot testing.
Method Long Short Overall
J F J&F J F J&F J F J&F
UniRef++ [79] (zs) 14.1 10.8 12.5 9.0 8.2 8.6 11.6 9.5 10.5
UniRef++ (ft) 19.2 15.1 17.2 12.3 11.7 12.0 15.8 13.4 14.6
Table 15: Region caption performance on RefCOCOg dataset.
Method Sa2VA-4B OMG-LLaVA [99] Osprey [92] GLaMM [66] GRIT [81] Kosmos-2 [64]
METEOR 17.3 15.3 16.6 16.2 15.2 14.1
Comparison with recent Video MLLM models. Recently, several works (including the not published work) also combine MLLMs into video referring segmentation tasks. As shown in Tab. 11, we list and compare these methods for fair comparison on three datasets. In particular, we also include recent work, VideoGLaMM, and a combined strong baseline (SAM2 + GlaMM). The combined baseline requires more computation and parameter costs than our model. As shown in that table, our model still achieves the best results with a 4B model size, along with the ability for image/video chat, GCG, and visual prompt understanding. To be noted our model achieves much stronger results than the combined baseline (SAM2 + GlaMM); this further shows the effectiveness of our joint-co-training and the necessity for task unification.
Comparison with vision expert models. In Tab. 12, we further compare the recent vision expert models, which are specifically designed for referring segmentation. Note that we list the best model results. Our best model with 26B mode size still achieves stronger results among these vision expert models.
Effectiveness of training dataset on more methods. We further show the effectiveness of our training dataset on more methods. In particular, we choose the previous strong vision expert, UniRef++. As shown in Tab. 13, with our training set, we find a significant improvement. This indicates that our automatic data engine narrows the domain gaps and improves the performance of existing Ref-VOS models.
Sa2VA with more MLLMs. In Tab. 13, we further include our Sa2VA with other MLLMs. We choose QWen2-VL. Compared with InternVL2 (the default model in our paper), we find a little performance drop in referring segmentation tasks. We will explore more MLLMs in future work.
1 Input: Video length $N$; Number of key frames $M$; Video frames $S_{N}$ ($X_{1}$, $X_{2}$, $X_{3}$$\ldots$, $X_{N}$); Language description $T$;
2 Output: Sequence of masks $M_{1}$, $M_{2}$, $M_{3}$$\ldots$, $M_{N}$;
3 Run: Sa2VA Model for Ref-VOS;
4 Extract key frames: $S_{M}$leftarrow $S_{N}$;
5 Visual embeddings: $E_{v}$leftarrow Encoder($S_{M}$);
6 Language embeddings: $E_{l}$leftarrow Encoder($T$);
7 Answers: $A$leftarrow LLM($\{E_{v},E_{l}\}$);
8 Prompt embedding: $P_{l}$leftarrow Linear(Find($A$, ’[SEG]’));
9 for $i=1,2,\ldots,M$ do
10       SAM-2 feature: $F_{i}$leftarrow Encoder($X_{0}$);
11       Mask: $M_{i}$leftarrow Decoder($\{P_{l},F_{i}\}$);
12       Update Memory: $Mem$leftarrow Cross-Attention($\{Mem,M_{i}\}$);
13
14for $i=M+1,M+2,\ldots,N$ do
15       SAM-2 feature: $F_{i}$leftarrow Encoder($X_{0}$);
16       Mask: $M_{i}$leftarrow Decoder($\{Mem,F_{i}\}$);
17       Update Memory: $Mem$leftarrow Cross-Attention($\{Mem,M_{i}\}$);
18
emit $M_{1}$, $M_{2}$, $M_{3}$$\ldots$, $M_{N}$;
Algorithm 1 Ref-VOS Inference Pipeline
Results on visual prompts understanding tasks. We also report visual prompt understanding tasks, following previous works [92, 99]. As shown in Tab. 15, our method also achieves the best results among recent visual prompt understanding models. This indicates that Sa2VA can also generate strong region-aware captions.
Appendix B Model Details and Implementation Details
Sa2VA model details of InternVL-2 baseline. In InternVL-2, for each video frame, we first resize it to 448 $\times$ 448, and then compress it into 256 tokens using the vision encoder of InternVL-2. The five key frames are arranged in temporal order and placed before the question. We modified the tokenizer and added the “[SEG]” token. The hidden state of the last layer in the LLM corresponding to the “[SEG]” token is used to generate the language prompt. During the inference process, if “[SEG]” is not present, it is assumed that the described object is not present in the video. If there is only a single image input, we follow InternVL-2’s dynamic resolution approach [9] to capture detailed information from the image.
Sa2VA model details of Qwen2-VL baseline. In Qwen2-VL [74], for each video frame, we first resize it to 448 $\times$ 448, and then compress it into 256 tokens following InternVL-2 for fair comparison. The five key frames are arranged in temporal order and placed before the question. We use M-RoPE as in Qwen2-VL [74]. We modified the tokenizer and added the “[SEG]” token. The hidden state of the last layer in the LLM corresponding to the “[SEG]” token is used to generate the language prompt. During the inference process, if “[SEG]” is not present, it is assumed that the described object is not present in the video. If there is only a single image input, use the same dynamic resolution approach as in InternVL-2 for fair comparison.
Sa2VA model details of Visual Prompt Understanding. We implement mask-based visual prompt understanding simply and effectively by cropping image features. Specifically, given an object mask, we resize it to match the image feature size and extract and flatten the image features from the overall view corresponding to the mask. The object will be represented by these vision tokens belonging to it. The format is: “Object1: <vp><IMG_FEAT>, …, <IMG_FEAT></vp>." “<vp>" and “</vp>" are two special text tokens that indicate the beginning and end of visual prompt features. In subsequent dialogue, this object will be referred to as “object1". To obtain the ability to understand the visual prompt, we jointly co-train the model with Osprey-727K [92] and the datasets used in the main paper.
Table 16: Comparison with previous Ref-VOS benchmarks.
Property DAVIS17-RVOS ReVOS Ref-YT-VOS MeVIS Ours
Short Text
Long Text
Large Object Motion
Large Camera Motion
Heavy Occlusion
Table 17: Statistics comparison with other Ref-VOS datasets.
Datasets Video Object Expression Mask Avg. Len.
Ref-YTVOS 3,978 7,451 15,009 131k 9.68
MeViS 2,006 8,171 28,570 443k 7.07
ReVOS 1,042 9,084 35,074 273k 10.48
Ref-SAV 37,311 72,509 72,509 6.0m 83.6
Ref-VOS inference pipeline with SAM-2 knowledge. As described in Algorithm 1, the pipeline begins by extracting the first five frames of the input video as keyframes, ensuring that they capture the critical context for the following processing. These key frames are fed into CLIP and flattened to visual sequential tokens for LLM processing. The LLM takes the visual and language tokens as input and uses these tokens to extract information about the video to generate the “[SEG]” token. In SAM-2, the prompt encoder encodes boxes or clicks to prompt embeddings for object referring. Different from SAM-2, we use two linear layers to project the “[SEG]” token into the language prompt embedding, which serves as an extension of the SAM-2 prompt encoders. With the language prompt embedding, we use the SAM-2 decoder to generate the masks of the key frames. Then, we use the memory encoder of SAM-2 to generate a memory based on the output key-frame masks. Finally, the memory attention in SAM-2 generates the remaining masks using the memory generated from the key-frame and previous non-key-frame masks.
Appendix C More Details on Ref-SAV Dataset
Comparison with existing Ref-VOS benchmarks. In Tab. 16, we further compare the previous existing Ref-VOS benchmarks from five different aspects: short text, long text, large object motion, large camera motion, and heavy occlusion. As in Table 17, our dataset has more detailed captions (Avg. Len. refers to the average length of the referring description) and many more masks (about 10x). The proposed data can lead to better instruction-following capability for complex descriptions. The previous benchmarks only contain partial aspects of these five challenging cases. On the other hand, our benchmark is built from SAM-2 [67] with long text, which is more challenging than previous benchmarks. This is why the previous Ref-VOS models cannot achieve good results on our benchmark.
Appendix D Visualization Results
Figure 5: Visualization results on image referring segmentation task.
Results on image referring segmentation. As shown in Fig. 5, we visualize the image referring segmentation task. With different language descriptions, our Sa2VA can segment different masks with the clues from the description.
Figure 6: Visualization results on video referring segmentation.
Results on video referring segmentation. As shown in Fig. 6, our method performs well in handling diverse and challenging conditions, showcasing a remarkable ability to adjust and perform effectively even in occlusion scenes or highly dynamic environments.
Figure 7: Visualization results on visual prompt understanding task. We use the masks predicted by our model under the GCG task as visual prompts, and generated region-level descriptions for these masks. The object masks and their captions for the corresponding region are highlighted in the same color.
Results on visual prompt understanding. As shown in Fig. 7, our method can generate descriptions based on the visual prompts. The descriptions generated by our method demonstrate a high degree of contextual awareness, effectively capturing details within the visual cues.
Figure 8: Visualization results on GCG tasks. Top: our method. Bottom: OMG-LLaVA [99]. Note that, our method has stronger and fined-grained grounding ability and text alignment than OMG-LLaVA [99], previous strong baseline.
Results on GCG task. In Fig. 8, we further show several visual examples in grounded caption generation tasks. Compared with the previous SOTA model, OMG-LLaVA, our model achieves better results in both mask quality and text-mask alignments. The former indicates better-aligned segmentation outputs, while the latter shows the latter indicates good text-to-region alignments.
Figure 9: Visualization failure cases.
Appendix E Failure Cases and Future Work
Failure cases of Sa2VA and Future work. Despite Sa2VA achieving stronger results on various ref-segmentation and video ref-segmentation datasets. However, there are still several cases of failure that need to be explored. We identify two shortcomings of Sa2VA in our future work. One is a long video with hard, distinguished referring examples, as shown in Fig 9. This is because our Sa2VA works in online mode without knowing the entire video content to align the long, complex text. The other is how to improve the VQA tasks (in image and video) without hurting the referring segmentation tasks. As shown in the ablation table of the main paper, increasing the scale of VAQ data leads to a performance drop in referring segmentation tasks. Thus, how to balance both types and keep the grounded prediction knowledge is still unsolved.
Future directions on Ref-SAV benchmarks. As discussed in Tab. 16, long text, occlusion inputs, and longer video with camera motion are critical problems in our benchmarks. Thus, the new models need to design a more robust Ref-VOS system to handle these challenges with several specific designs, including more robust memory designs. In addition, understanding long text and improving long text grounding ability is also important to explore.
Board impact. Our research explores a unified solution of grounded understanding, VQA, and visual prompt understanding. Our work has application potential in the fine-grained promptable video understanding the task, in particular, in zero-shot inference mode, as shown in our teaser in the main paper. In addition, with solid experiment results, our exploration will inspire the development of interactive video-dense understanding in the MLLM community.